Recently, Steve replaced the HeNe which was sourcing the BS & PRM OL. After replacement, no one checked the beam sizes and we've been living with a mostly broken BS OL. The beam spot on the QPD was so tiny that we were seeing the 'beam is nearly the size of the segment gap' effect.
Today I removed 2 of the lenses which were in the beam path: one removed from the common PRM/BS path, and one removed from the PRM path. The beams on both the BS & PRM got bigger. The BS beam is bigger by a factor of 7. I've increased the loop gains by a factor of 6 and now the UGFs are ~6 Hz. The loop gains were much too high with the small beam spots that Steve had left there. I would prefer for the beams to be ~1.5-2x smaller than they are now, but its not terrible.
Many of the mounts on the table are low quality and not constructed stably. One of the PRM turning mirror mounts twisted all the way around when I tried to align it. This table needs some help this summer.
In the future: never try locking after an OL laser change. Always redo the telescope and alignment and check the servo shape before the OL job is done.
Also, I reduced the height of the RG3.3 in the OL loops from 30 to 18 dB. The BS OL loops were conditionally stable before and thats a no-no. It makes it oscillate if it saturates.
|